# Fuzzy Ontology Embeddings and Visual Query Building for Ontology Exploration

### v2

@@ 5.3.3 Visualization Controls
The left region of FuzzyVis contains controls and panels for manipulating the primary visualization (Figure 2.C). To illustrate, panels for the treemap visualization are shown in Figure 5 This region holds three components. First, the highlight panel, for setting rules to stain the primary visualization with colors (Figure 5(a)). Second, the visualization control panel, for manipulating the behavior of the primary visualization (Figure 5(b)). Last, is a help panel to explain to unfamiliar users how to interact with the system. The left region’s purpose is to allow users to tune the visualization to more closely meet their needs. As users need vary, a single set of static visualization parameters is often insufficient. For example, a user with a smaller monitor may require a more horizontal tiling for a treemap in order to be able to read the labels without mousing over the cells.
@@
The highlight panel (Figure 5(a)) allows users to annotate the primary visualization. In this panel, users construct rules and assign colors to highlight ontology concepts. Concepts which if they satisfy multiple rules will mix the highlight colors. Users can filters based on concept label, subtree size, number of children, or other criteria when making rules. Rules that restrict on string values can either require an exact match or just substring matching. Rules that restrict on numerical values can set minimums, maximums, equality, and ranges. Additionally, users can negate rules or toggle them on and off.
@@
For example, a highlighting rule may specify: color blue all concepts with labels that contain bulbar (the first rule in Figure 5(a)). This rule stains concepts such as Bulbar Palsy and Pseudobulbar signs blue making them significantly more noticeable. Highlighting allows users to quickly identify concepts of interest to form landmark and neighborhood knowledge. Isolated colored concepts exhibit uniqueness, while groups of similarly colored concepts could bear similarity to each other depending on the rules. Furthermore, users can set in the controls to only display highlighted concepts in the primary visualization. For large ontologies with many concepts some visualizations will densely pack concepts making it hard to distinguish them (ex. treemaps, or network graphs). In such scenarios, highlighting provides needed visual marks for distinguishing concepts. Moreover, this panel’s function can be expanded to support ontology-specific rules, depending on an ontology’s properties.
The visualization control panel is used to configure the behavior of the primary visualization. Here users can adjust the parameters that determine how the primary visualization is displayed. Depending on the selected visualization options will differ. For example, Figure 5(b) is the controls for a nested treemap visualization. For such treemaps, settings include:
@@
Depth. The max displayed depth from the selected concept. Higher values will densely pack the interior of cells and allow quick traversal to lower layers. Where as, lower values improve clarity.
Tiling Ratio. The shape of the treemap cells. User preference for more or less horizontal and vertical cells is controlled with this setting. Depending on screen size and ontology content some ratios can improve or reduce clarity significantly. For example, a horizontal ratio may make long labels easier to read, but with too many children concepts there many not be enough vertical space to show the label when all concepts are stacked as plates.
Scaling. The value that determines the size of concepts. Users interested in metadata such as subtree size or number of children can choose to scale concepts based on these values. A very large concept contains the majority of the ontology and can be considered more important. On the other hand, equalizing scaling lets users see concepts that would be hidden due to the dominance of others.
Visibility. Which concepts are displayed. Based on highlighting rules determine which concepts are shown in the primary visualization. Users only interested in specific concepts can hide those they deem irrelevant.
Normalization. Adjust the scaling so that minor concepts are more visible and dominate ones take up less space. Another means by which users can control how concepts are displayed.
@@
Focus Mode. Enable mouse over distortion of the visualization. Users can dynamically adjust the scaling of concepts to focus on what interests them regardless of the ontology’s structure.
Showing Layer Numbers. Annotate concepts with their layer number. Some users may struggle to distinguish depth in some visualizations. For example, in a network graph it can be difficult to distinguish how deep a concept is in a given subtree.
@@
Toggling similarity stain. Stain the concepts with colors based on the query results. Users exploring queries can quickly see what parts of the ontology correspond to their query.
@@
Some of these settings are universal, others, are visualization specific. In general, the purpose of these controls is to align the visualization to users’ needs. As ontologies can have a great amount of variety, default settings are not guaranteed to be optimal—allowing users to adjust the settings obviates this issue.
The last panel, help and keybindings provides users with information on how the system functions. Users can hide this view if it is unneeded. Though a full interactive tutorial would be more optimal it is beyond the scope of this prototype.
@@ 5.3.4 Concept Panels
The first of the right region of FuzzyVis are the concept panels (Figure 2.D). These panels have two primary functions: organize a collection of concepts for future use and provide concept details. As users explore ontologies, they encounter concepts of interest that they may want to revisit or use in queries. These concepts may be located in disparate regions making them inconvenient to access. As well, users need concept details to evaluate relevance and suitability. In HPO, users investigating skull-related conditions may be interested in the Abnormal cranial nerve physiology and Abnormality of the head concepts. Yet, neither of these concepts are conveniently close to each other. Nor without seeing details can users confirm that these concepts truly meet their requirements. When users store these concepts in the collection they can quickly navigate to and between them. Overall, the concept panels are present to aid users in their navigation and query formulation.
The collection panel (Figure 6(a)) stores concepts of interest for easy access. Users drag and drop, use any of the collection keybindings, or pin concepts to add them to the panel. Additionally, users control which concepts are pinned here. The primary visualization annotates with a visual mark any pinned concepts for easier identification. The concepts within the collection can be dragged to any other component that takes concepts as inputs. The highlight component (Figure 5(a)) constructs a rule to match dropped concepts. The primary visualization refocuses onto the dropped concept. As well, when dropped on, any query being built will have the concept added to it. Users’ collections allow them to quickly navigate back to concepts of interest and not have to remember the full names of every concept that they are interested in. A user selecting Abnormality of the head can relocate the primary visualization there regardless of where they currently are exploring.
The details panel (Figure 6(b)) shows information about the currently selected concept. Dependent on the ontology there may be more or fewer entries. However, common among most ontologies are values for: label, IRI, parent, and children (concepts which possess a subclass relationship with a given concept). Thus, this panel can show a significant amount of critical information. For example, HPO includes concept definitions which provide clarity as to what a concept represents. While the primary visualization provides high-level spatial information for the ontology concepts, it forgoes details in favor of visual clarity. As such, the details panel is necessary to provide low-level detailed information to users for individual concepts.
@@ 5.3.5 Query Building Panels
The second section of FuzzyVis ’s right region are the query building panels (Figure 2.E). Users interact with these components to create, resolve, and analyze queries. Hierarchical exploration of ontologies exposes users to only a portion of the semantics and structure. Furthermore, analyzing individual concepts one at a time is inefficient. Thus, as users explore they may want to verify their inferences or locate unknown, yet identifiable, concepts. The query building panels enable users to perform such non-linear exploration and develop an understanding of the structure and semantics of ontologies. We discuss this in-depth in our usage scenario (Section 5.4).
Users construct their queries with the query builder control (Figure 7(a)). Queries are composed of concepts, other queries, and support standard logical operators (AND, OR, and NOT). These operators correspond to intersection, union, and negation respectfully. When users finalize their queries, FuzzyVis ’s front-end passes them to the back-end query resolver which converts them to composite concepts. These composite concepts have their embeddings compared to the ontology’s primitive concepts before a list of top-k most similar concepts are returns. There is no restriction on which concepts can be added to a query, nor is there anything stopping a user from creating a logically inconsistent statement. Such a query is useless when resolved with traditional embeddings. However, our fuzzy ontology embeddings can still capture semantic information from some such queries.
The query results are shown in FuzzyVis ’s query results panel (Figure 7(b)). Here users can see the degree to which each primitive concept is similar to their query’s composite concept. Users can expand any of the results to show concept details. As well, users can drag the concepts from the results list and use them in future queries or in other regions of the system. The similarity stain option in the visualization control panel works with these results. When this option is toggled on the similarity values of the query results stain the primary visualization with a matching tone. The query resolver uses cosine similarity to determine the similarity score. Because FuzzyVis stores primitive concept embeddings in a vector database queries resolve nearly instantly.
@@ 5.4 Usage Scenario
The following usage scenario demonstrates the features of FuzzyVis and the practical application of fuzzy logic queries for ontology exploration. We use the HPO for this scenario. We assume the user knows basic medical terms but is unfamiliar with HPO. The embedding used is an -embedding (as described in Section 4.1) with . For fuzzy operators, we use the product t-norm for conjunction, its t-conorm for disjunction, and the standard negator for negation.
@@
A physician, henceforth referred to as the user, has a patient who has difficulty speaking and swallowing. This patient has no history of immune disorders and tests reveal no signs of infection. To help the patient, the user wants to accomplish two goals. One, they want to identify possible conditions that could explain the patient’s symptoms. Two, they wish to learn about any secondary ailments that the patient may experience now or in the future. The user looks to HPO as it contains information about abnormal conditions. However, the user needs a tool to assist with exploring the ontology due to its complexity and size. Thus, the user launches FuzzyVis to explore HPO in order to help their patient.
@@
The user opens FuzzyVis ’s front end in their browser of choice and loads HPO with the previously mentioned parameters. After initializing, FuzzyVis is ready to use and defaults the primary visualization to a nested treemap. The user sees the initial state of the primary visualization and adjusts the settings in the visualization control panel. First, the user sets the scaling of each cell to be equal as the relative size of concepts is not relevant to their goals. Second, the user sets the tiling so that the concepts clearly show their labels. Lastly, they set the depth to a low value as HPO has many terms and showing multiple layers is overwhelming. Now each concept is clearly visible to the user. Next, the user creates a few highlight rules for key terms they are interested in. For this example, the highlight rules are to stain any concept with the words voice, speech, and swallowing in their label for easier identification. After this set up the user can now quickly identify initial concepts of interest. Users of FuzzyVis can with little effort setup the primary visualization in ways that streamline their future searches.
Seeing the concept Phenotypic abnormality, the user selects it and begins their search. During their search, the user identifies several child concepts related to their patient’s condition. These concepts are pinned (for easier identification) and added to the concept collection for use in queries. To explore more specific concepts, the user activates focus mode and starts browsing the contents of Phenotypic abnormality’s children. They see that Abnormality of the voice does not contain what they are looking for; the user collects it for query use regardless. Additionally, the user spots Abnormality of the immune system and adds it to the collection so it can be negated in future queries. Hoping to quickly find something, the user interacts with the search bar and looks for concepts whose labels contain the word speech. Among the keyword search results is the concept Slurred speech, a direct match for the user. Selecting Slurred speech takes the user deeper into the ontology, where they then pin it. Wanting to know where in the ontology they are, the user toggles on the depth numbers and sees that Slurred speech is eight layers deep and has no children. Then, the user increases the depth to view the full ancestry and spots that Slurred speech is within the Abnormality of the nervous system subtree. Now that the user knows where to look, they select Abnormality of the nervous system, reduce the shown depth. and continue their search here. Already the user has started to identify landmark concepts related to their problem, has learned the routes between them, and is now aware of the neighborhood in the ontology that contains concepts of interest. FuzzyVis allows for fast and seamless exploration; users can quickly see concept contents at a glance and freely navigate ontologies.
@@
The user’s exploration continues for only a bit longer. Settling on Abnormality nervous system physiology the user encounters the concept Dysphagia within it. The focus mode alternative views and the concept details panel allow the user to quickly see the definition of this concept. As Dysphagia is the medical term for difficulty in swallowing, and with the prior discovery of Slurred speech, the user is now confident that they can start constructing queries. The user formulates their query as the presence of Slurred speech and Dysphagia while excluding Abnormality of the immune system. Using the query building panels the user drags concepts from the collection to the panel to create their query. Ultimately, the user defines the composite concept :
@@
The user sends in their query, FuzzyVis ’s back-end quickly processes it, and the front-end updates nearly instantly with results. Among these results are concepts such as Pseudobulbar paralysis, Pseudobulbar signs, and Abnormal esophagus physiology. Wanting to locate these concepts in the ontology, the user turns on the similarity stain and spots the Pseudobulbar signs concept. With focus mode on, the user locks and splits Pseudobulbar signs to show its details and expands it to show Pseudobulbar paralysis. The results of both the similarity stain and this focusing can be seen in Figure 8. The pseudobulbar concepts refer to neurological conditions that explain the patient’s symptoms. On the other hand, Abnormal esophagus physiology is an alternative direction for investigation as it can also explain the symptoms. These results have now narrowed down potential causes that the user can investigate to diagnose their patient. Finding these concepts with keyword search would have been difficult for the user as their vocabulary did not perfectly match the ontology’s vocabulary. Moreover, the user would not know how to formulate a full SPARQL query to find these concepts. Users of FuzzyVis do not need to write out queries, instead, they build them through interactions with the interface by dragging and dropping concepts and other queries to define complex concepts. FuzzyVis ’s visual query building obviates the need for users to precisely spell out concept labels and by supporting nesting queries helps users separate the components of their queries into understandable segments. In allowing users to create fuzzy queries, FuzzyVis lets users find concepts relevant to their needs even when their requests are imprecise.
An example of the imprecise kinds of queries possible is: our scenario user’s interest in finding secondary ailments, a task they did by updating their query to exclude Abnormality of the voice. This would be an illogical query. However, because the -embedding contains fuzzy membership across the ontology the user believes that this query may still offer useful insights. Thus, the user sends in the updated query :
@@
returns concepts related to digestive issues. This is expected as Dysphagia is part of the Abnormal esophagus physiology subtree which belongs to Abnormality of the digestive system. Thus, the condition of the patient may have consequences for the digestive system. The fuzzy ontology embeddings we use capture such relationships and continue to preserve them in complex composite concepts created by users. Armed with this new knowledge the user can ask their patient further questions to narrow down their condition. To check for Pseudobulbar signs they can inquire if the patient has found themselves more irritable and prone to emotional outbursts. To check for Abnormal esophagus physiology the user can simply inspect the patient’s throat.
@@
This usage scenario highlights how FuzzyVis ’s features and fuzzy ontology embeddings can be used to explore ontology content. Despite having no prior knowledge of HPO’s structure and content, or query languages like SPARQL, the user was able to find relevant concepts and develop an understanding of the ontology. The system’s interface—composed of interactive visualizations—empowers users to visually build queries and explore the ontology. Furthermore, the use of fuzzy logic enables users to create queries even without expert knowledge of the ontology. These features of FuzzyVis and fuzzy ontology embeddings represent a way of exploring ontological information that is more flexible and user-friendly.

### v3

@@ 5.3.3 Visualization Controls
The left region of FuzzyVis contains controls and panels for adjusting the primary visualization (Figure 2.C). To illustrate, Figure 5 shows the panels for a nested-treemap visualization. This region has three components: the highlight panel for applying color-based rules (Figure 5(a)), the visualization control panel for adjusting the primary visualization’s behavior (Figure 5(b)), and a help panel explaining system interaction. These controls let users tailor the visualization to their needs, as static parameters are often insufficient. For example, a user on a small monitor may require a more horizontal treemap layout to read labels without hovering.
@@
The highlight panel (Figure 5(a)) allows users to annotate the primary visualization by creating rules that assign colors to concepts. When a concept satisfies multiple rules, colors are blended. Rules can filter by label, subtree size, number of children, or other criteria; string-based filters may require exact or substring matches, while numerical filters can specify ranges, minimums, maximums, or equality. Rules can also be negated or toggled on and off.
@@
For example, a rule may color concepts blue if their labels contain bulbar (Figure 5(a), first rule), making concepts such as Bulbar Palsy and Pseudobulbar signs stand out. Highlighting helps users quickly identify concepts of interest to form landmark and neighborhood knowledge. Isolated colored concepts may indicate uniqueness, while clusters can suggest similarity. Users can also choose to display only highlighted concepts, which is especially useful in dense visualizations such as treemaps or network graphs. Highlighting thus provides essential visual cues for navigation and can be extended to support ontology-specific rules when needed.
The visualization control panel configures the behavior of the primary visualization, with available options varying by visualization type. For example, Figure 5(b) shows the controls for a nested-treemap, which include:
@@
Depth. Maximum displayed depth from the selected concept. Higher values pack more content for rapid traversal; lower values improve clarity.
Tiling Ratio. Controls the horizontal–vertical balance of treemap cells. Screen size and ontology structure affect which ratios best preserve readability; for example, a strongly horizontal layout aids long labels but may have insufficient vertical space when many concepts are present.
Scaling. Determines concept size. Scaling by subtree size or number of children emphasizes dominant concepts, while equal scaling reveals smaller ones that might otherwise be hidden.
Visibility. Controls which concepts are shown, often based on highlighting rules, allowing irrelevant concepts to be hidden.
Normalization. Adjusts scaling to make small concepts more visible and large ones less dominant.
@@
Focus Mode. Enables mouse over distortion, dynamically enlarging concepts of interest regardless of ontology structure.
Showing Layer Numbers. Annotates concepts with their depth, useful in visualizations (e.g., network graphs) where depth is hard to gauge.
@@
Toggling Similarity Stain. Colors concepts based on query results, quickly showing which areas of the ontology match.
@@
Some settings are universal, while others are visualization-specific. These controls let users adapt the visualization to their needs, as default settings are rarely optimal for all ontologies.
The final panel—help and keybindings—provides usage guidance and can be hidden if not needed. While a full interactive tutorial would be ideal, it is beyond the scope of this prototype.
@@ 5.3.4 Concept Panels
The first section of the right region of FuzzyVis contains the concept panels (Figure 2.D), which serve two purposes: organizing the collection of concepts for later use and displaying concept details. As users explore ontologies, they may find concepts worth revisiting or using in queries. These concepts can be scattered across the ontology, making them inconvenient to access without a collection. Details are also needed to confirm relevance. For example, in HPO, a user studying skull-related conditions may collect both Abnormal cranial nerve physiology and Abnormality of the head, despite their distance in the hierarchy. Storing them in the collection allows quick navigation between them. Overall, the concept panels aids in navigation and query formulation.
The collection panel (Figure 6(a)) stores user-selected concepts for easy access. Users can add concepts to the panel by drag-and-drop, by using collection keybindings, or by pinning them. Additionally, users can control which concepts are pinned here. Concepts in the collection can be dragged to other components, such as the highlight panel, to create a matching rule; the primary visualization, to refocus; or the query builder, to add them to a query. Collections save users from having to remember full concept names and allow rapid return to points of interest—for instance, quickly relocating the view to Abnormality of the head.
The details panel (Figure 6(b)) displays information for the currently selected concept. Depending on the ontology, the number of attributes varies. Most ontologies’ concepts include labels, IRIs, parents, and children. In HPO, for example, concept definitions add clarity to a term’s meaning. While the primary visualization provides high-level spatial information, it forgoes details in favor of visual clarity. As such, the details panel delivers essential low-level information for individual concepts.
@@ 5.3.5 Query Building Panels
The second section of the right region contains the query building panels (Figure 2.E), which allow users to create, resolve, and analyze queries. Navigating an ontology hierarchically only reveals part of its structure and semantics, and inspecting concepts one at a time can be inefficient. Query building enables non-linear exploration, helping users verify inferences or discover related concepts. We discuss this in-depth in our usage scenario (Section 5.4).
Queries are created in the query builder control (Figure 7(a)) from concepts, other queries, and standard logical operators (AND, OR, NOT), corresponding to intersection, union, and negation. Once finalized, the front-end sends the query to the back-end resolver, which converts it to a composite concept. This composite concept’s embedding is compared with primitive concept embeddings, returning the top- most similar concepts. Queries may contain any concepts, including logically inconsistent combinations—which, while ineffective when resolved with traditional embeddings, can still yield meaningful results with our fuzzy ontology embeddings.
Results are displayed in the query results panel (Figure 7(b)), showing similarity scores for each primitive concept. Users can expand results for details, or drag them into new queries or other components. The similarity stain option in the visualization control panel colors the primary visualization according to these similarity values. The query resolver uses cosine similarity, and because primitive concept embeddings are stored in a vector database, queries resolve nearly instantly.
@@ 5.4 Usage Scenario
The following scenario illustrates how FuzzyVis supports ontology exploration and the practical applications of fuzzy logic queries. We use the HPO for this scenario. We assume the user knows basic medical terms but is unfamiliar with HPO. The embedding used is an -embedding (as described in Section 4.1) with . For fuzzy operators, we use the product t-norm for conjunction, its t-conorm for disjunction, and the standard negator for negation.
@@
A physician, henceforth the user, has a patient who has difficulty speaking and swallowing, with no history of immune disorders, and no signs of infection. To help this patient, the user has two goals. One, they want to identify possible conditions that could explain the symptoms. Two, they wish to learn about any secondary ailments that the patient may experience now or in the future. The user knows that HPO contains information about abnormal conditions, but its size and complexity necessitates the use of an exploratory tool. Thus, the user launches FuzzyVis to explore HPO and help their patient.
@@
After FuzzyVis loads HPO, its front-end—displaying a nested-treemap as the primary visualization—presents the ontology to the user. The user starts their exploration by adjusting the settings in the visualization control panel. First, the user sets cell scaling to be equal, as the relative size of concepts is irrelevant to the their goals. Second, the user sets the tiling so that labels are clearly visible. Lastly, they set visible depth to a low value—HPO has many terms and seeing too many layers is overwhelming. Afterwards, the user creates highlight rules to stain concepts containing voice, speech, or swallowing for quick identification. With the user’s initial setup complete, they can now quickly identify concepts of interest. Users of FuzzyVis can, with little effort, set up the primary visualization in ways that streamline their future searches.
Seeing the concept Phenotypic abnormality, the user selects it and begins their search. During their search, the user encounters concepts related to their patient’s condition. These concepts are pinned (for easier identification) and added to the concept collection for use in queries. To explore more specific concepts, the user activates focus mode to drill into Phenotypic abnormality’s children. They see Abnormality of the voice, and although it does not contain what they are looking for, it is collected for query use regardless. Additionally, the user spots Abnormality of the immune system and adds it to the collection, to be negated in future queries.
Wanting to quickly find something, the user inputs speech into the search bar. Among the results is the concept Slurred speech, which matches the patient’s symptoms. Selecting it takes the user deeper into the ontology, where they then pin it. Unsure of where they are, the user toggles on the depth numbers and sees that Slurred speech is eight layers deep and has no children. Then, the user increases the depth to view the full ancestry and spots that Slurred speech is within the Abnormality of the nervous system subtree. The user selects this broader concept, reduces depth for clarity, and continues exploring.
Already the user has started to identify landmark concepts related to their problem, has learned the routes between them, and is now aware of the neighborhood in the ontology that contains concepts of interest. By quickly identifying this information, the user efficiently narrows the search space. FuzzyVis’s support for rapid, seamless navigation allows users to quickly see concept contents at a glance and freely navigate large ontologies.
@@
The user’s exploration continues until, within Abnormality of the nervous system child concept Abnormal nervous system physiology, they encounter Dysphagia. The focus mode alternative views and concept details panel show the user that Dysphagia is the medical term for difficulty swallowing. Combined with the earlier discovery of Slurred speech, the user feels ready to construct queries. The user formulates their query as the presence of Slurred speech and Dysphagia while excluding Abnormality of the immune system. They do this by dragging concepts from the collection to the query building panel to form the composite concept :
The user submits their query, FuzzyVis’s back-end instantly resolves it, and the front-end updates to show the results. Among these results are concepts such as Pseudobulbar paralysis, Pseudobulbar signs, and Abnormal esophagus physiology. Wanting to locate these concepts, the user enables the similarity stain and spots the highlighted Pseudobulbar signs. Using the focus mode, the user locks and splits Pseudobulbar signs to expand it and see its details. Notably, this reveals Pseudobulbar paralysis within Pseudobulbar signs. Figure 8 shows the combined effect of this similarity stain and focus mode, highlighting the pseudobulbar concepts. These pseudobulbar concepts point to neurological causes that explain the patient’s symptoms. On the other hand, Abnormal esophagus physiology offers an alternative explanation for the symptoms.
@@
These results have now narrowed down potential causes that the user can investigate to diagnose their patient. Such concepts would have been difficult to find with keyword search, as the user’s vocabulary does not match the ontology’s precisely. Moreover, the user would have been unable to find these concepts with SPARQL queries for this same reason. With FuzzyVis, users build queries visually by dragging and dropping concepts and other queries, without needing to type exact labels. Furthermore, nesting queries helps users separate the components of their queries into understandable segments and construct complex concepts. By supporting fuzzy queries, FuzzyVis enables users to find relevant concepts even when their requests are imprecise.
An example of an imprecise query is the user’s search for secondary ailments. The user creates this query by updating their original one to exclude Abnormality of the voice. In a non-fuzzy system, excluding Abnormality of the voice while searching for conditions related to speech and swallowing might seem counterintuitive or overly restrictive. However, because the -embedding captures fuzzy membership across the ontology, the query can still yield useful insights. The new query narrows the focus to concepts returned by the original that are unrelated to Abnormality of the voice. The updated query is:
@@
Submitting returns concepts related to digestive issues, which is expected since Dysphagia belongs to the Abnormal esophagus physiology subtree under Abnormality of the digestive system. This suggests the patient’s condition may affect the digestive system. The fuzzy ontology embeddings preserve such relationships even in complex composite concepts. Armed with this knowledge, the user can further question their patient to narrow down their condition. The user asks about irritability and emotional outbursts to check for Pseudobulbar signs, and inspecting the throat to check for Abnormal esophagus physiology.
@@
This scenario demonstrates how FuzzyVis ’s features and fuzzy ontology embeddings enable effective ontology exploration. Without prior knowledge of HPO’s structure or query languages like SPARQL, the user was able to find relevant concepts and understand their relationships. The interface’s interactive visualizations allows the user to build queries visually, while fuzzy logic supports the creation of useful queries even without expert knowledge. Together, these capabilities present a flexible and user-friendly approach to exploring ontological information.